OpenStudy (anonymous):

(X) |1| T (y) = |z| (Z) |y| Determine whether the function T: R^3->R^3 is a linear transformation

OpenStudy (freckles):

we need the following to hold: \[\text{ Let } v_1,v_2 \in \mathbb{R}^3\\ T(v_1+v_2)=T(v_1)+T(v_2) \\ \text{ and } \\ T(av)=aT(v) \text{ where } a \text{ is a scalar }\]

OpenStudy (freckles):

are you there?